# Tuning-Free Alignment of Diffusion Models with Direct Noise Optimization

### v1

@@ 3 Out-Of-Distribution Reward-Hacking in Noise Optimization
It has been observed that when aligning generative models with reward functions, they can be subject to reward-hacking, i.e., the optimized samples yield high rewards but do not possess the desirable properties. Generally, there are two different types of reward-hacking. The first is because the reward model is undertrained, which results in optimized samples scoring high rewards but remaining barely distinguishable from the samples of the pretrained distribution (Miao et al., 2024; Chen et al., 2024a). The second type relates to the generative model used – the optimized samples no longer fall within the support of the pretrained distribution after optimization. We denote this second type of reward-hacking as Out-Of-Distribution (OOD) Reward-Hacking. In this work, we will focus on this second type and reveal that OOD reward-hacking is a common issue in DNO.
We begin by presenting two examples using both a simple diffusion model from the toy example and the open-source image diffusion model SD v1.5 (Rombach et al., 2022). For the toy example, we revisit the pretrained distribution displayed in Figure 2, modifying the reward function to $r(x)=-(x[1]-1.4)^{2}-(x[2]-1.4)^{2}$. Figure 4(a) exhibits the pretrained distribution of the diffusion models, where we note that every sample stays within the support. However, as illustrated in Figure 4(b), after 1000 gradient steps of optimization for the reward function, all the generated samples become out-of-distribution. In our second example, we examine optimization for the "whiteness" reward—specifically, the average value of all pixels in an image—using SD v1.5 as the image diffusion model. We start with the prompt "black duck", with the initial image depicted in Figure 4(c). After 50 gradient steps of optimization for the whiteness reward, it becomes apparent that the generated samples diverge from the original prompt "black duck" and transform into a white duck, as evidenced in Figure 4(d).
@@
Cause of OOD Samples. In this work, one of our key contributions is to identify one critical cause of OOD reward-hacking in noise optimization. That is, the optimized noise vectors stray towards the low-probability regions of the high-dimensional standard Gaussian distribution; in other words, there is an extremely low chance of such noise vectors being sampled from the Gaussian distribution. As diffusion models are originally trained with Gaussian noise, when the noise vectors originate from these low-probability areas—such as vectors comprised entirely of zeros—the neural network within the diffusion models may incur significant approximation errors for these particular inputs. This error, in turn, prompts the generation of out-of-distribution samples. In the subsequent section, we introduce an innovative method to measure the extent to which noise is part of the low-probability region by leveraging the classical concentration inequalities for high-dimensional Gaussian distributions.
@@ 3.1 Quantifying Low-Probability Region via Concentration Inequalities
An important thing to note is that the standard Gaussian distribution is invariant to permutation, i.e., for any permutation matrix $\Pi$, if $z$ follows a standard Gaussian distribution, the permuted vector $\Pi z$ will have the same probability behavior. With this insight, to increase the robustness of the probability measure $p_{1}$ and $p_{2}$, a natural idea is to examine the probability of many permuted vectors. Specifically, given $q$ permutation matrices $\Pi_{1},...,\Pi_{q}$, we define the following indicator metric,
@@
Interpretation of $P(z)$. If the probability $P(z)$ is low, it implies that there exists a permutation matrix $\Pi_{i}$ such that the noise vector $\Pi_{i}z$ is in the low-probability region of the standard Gaussian distribution. Therefore, due to the permutation-invariant property, the noise vector $z$ is also less likely to be sampled from the standard Gaussian distribution. In practice, we found that setting $q=100$ and using randomly generated permutation matrices is an empirically good choice. In Appendix C, we provide some visualized empirical evidence to show that $P(z)$ serves as a good indicator for determining if the generated samples are OOD.
@@ 4 Tackling Non-Differentiable Reward Functions
As another major contribution of this work, we explore adapting the noise optimization approach to handle the optimization of non-differentiable reward functions by estimating the gradient with function values. Specifically, we explore three methods under this setting.
@@
Method 2. It worth to note that for the mapping $r\circ M_{\theta}(\cdot)$, only the gradient of the reward function $r(x)$ is not available, and we are still able to compute the gradient of $M_{\theta}(z)$. Therefore, a straightforward idea is to adopt a hybrid gradient approach—only to estimate the gradient of $r$, while using the ground truth gradient for $M_{\theta}(z)$. Specifically, we denote that the initial noise is $z$, and the generated sample is $x=M_{\theta}(z)$. Firstly, we can estimate the gradient of $\nabla r(x)$ in a similar fashion with the ZO-SGD (Nesterov and Spokoiny, 2017):
@@
where $\mu$ is the coefficient for perturbation, and $C_{1}$ is some constant. With the estimated gradient $H_{1}(x)$, we can use the following gradient $H_{1}(x)\cdot\nabla_{z}M_{\theta}(z)$ for optimization, where the main idea is to replace the ground truth $\nabla r(x)$ in the chain-rule of differentiating $r(M_{\theta}(z))$. We refer this method as Hybrid-1 in the following sections.
@@
Following a similar proof in (Nesterov and Spokoiny, 2017), we can also show that $H_{2}(x)\approx C_{2}\nabla r(x)$ for some constants $C_{2}$. As we can see, when computing the gradient (14), we ensure that we query the reward model $r(\cdot)$ with only samples that are within the manifold of the pretrained distribution. We refer this method as Hybrid-2. As we will see in Section 5.3, this Hybrid-2 method is significantly faster than the other two in terms of optimization speed. In Appendix D, we describe an efficient way to implement (13) and (14) with auto-differentiation technique.
@@ 5 Experiments
Figure 5: Upper row: Optimizing whiteness reward. Lower row: Optimizing blackness reward.
In this section, we aim to demonstrate the effectiveness of the method proposed above. For all subsequent experiments, we utilize Stable Diffusion v1.5 (Rombach et al., 2022) as the base model for noise optimization. For each figure, we perform the optimization using 1,000 different random seeds and report the average value along with the standard deviation (std) of the results. For comprehensive details regarding the implementation of our proposed methods, as well as information on hyperparameters, we refer readers to Appendices D and E. Additionally, we provide examples to visualize the optimization process in Appendix A.
@@ 5.1 Examining the Effectiveness of Probability Regularization
Setting. In this section, we design a novel experiment to demonstrate the effectiveness of probability regularization proposed in Section 3.2. We consider two settings: the first involves optimizing the whiteness reward, i.e., the average value of all pixels—the larger this value is, the whiter the image becomes—with the prompt "black <animal>", where the token <animal> is randomly selected from a list of animals. The second setting involves optimizing the blackness reward, i.e., the negative of the whiteness reward, with the prompt "white <animal>". The primary insight behind designing such experiments is the clear contradiction between the prompt and the reward, making it easier to trigger the OOD reward-hacking phenomenon. Moreover, it is straightforward to verify whether the generated samples fall within the support of the prompt by simply examining the color of the generated animals. In these experiments, we will compare the noise optimization process with and without probability regularization to see whether probability regularization can prevent the OOD reward-hacking phenomenon.
@@
Figure 6: Examples of optimized images with and without regularization.
Measuring the Degree of OOD. As observed in this experiment, the reward function contradicts the input prompt, leading to inconsistency between the generated samples and the prompt. Therefore, we utilize the CS (Radford et al., 2021), a commonly used metric for measuring the semantic similarity between images and text descriptions, to gauge the degree of OOD for the generated samples.
Results. In Figure 5, we first observe that adding the regularization term leads to a mildly slower optimization process. However, the generated samples are much more consistent with the prompt throughout the entire optimization process, as reflected by the CS curve. The trajectory of $P(z)$ further corroborates that it is a good indicator of the OOD phenomenon, as it is positively associated with the CS. It is worth noting that the regularization effect provided by probability cannot be fully replicated by simply employing early-stopping in non-regularized optimization, as illustrated in Figure 5. We offer qualitative examples in Figure 6 to highlight the differences between optimized samples with and without regularization. Figure 6(a) and 6(b) are optimized images with whiteness reward and prompt "black rabbit". Figure 6(c) and 6(d) are optimized images with blackness reward and prompt "white chicken". In the two pairs of images, we select the optimized images that achieved the almost same reward value from the optimizations with and without regularization, respectively. It is evident that at an equivalent level of reward, the non-regularized samples are more prone to being out-of-distribution, as the color of the generated animals is less likely to match the prompt.
Figure 7: A, H, P are short for Aesthetic Score, HPS Score, and Pick Score, respectively. The name for each line comprises the used reward model and whether the regularization is used. For example, A + w/ reg means optimizing aesthetic score with regularization. The $x$-axis is the gradient steps.
@@ 5.2 Benchmarking Three Reward Models Trained on Human Feedback Data
Setting. In this section, we investigate the performance of the proposed method using three common reward models trained from human feedback data, specifically Aesthetic Score (Schuhmann et al., 2022b), HPS-v2 score (Wu et al., 2023), and PickScore (Kirstain et al., 2023), respectively. In this experiment, we also compare noise optimization with and without probability regularization. However, compared to the reward function used in the previous section, using these reward models present a lesser chance for the optimized image to be OOD. In this case, to measure the benefit of probability regularization in maintaining the quality of the generated sample, we consider using the other two reward models as test metrics when optimizing one of them. For example, when optimizing the Aesthetic Score, we use the HPS Score and Pick Score as test metrics.
@@
Comparing to Existing Alignment Methods. We also summarize the performance of our proposed method in Table 1 and compare it to the major existing alignment methods discussed in the introduction. As we can see, the proposed method can easily outperform state-of-the-art tuning-based alignment methods without any fine-tuning on the network models, within a reasonable time budget for generation. We found that another tuning-free method, LGD (Song et al., 2023a), performs badly with these complex reward functions, since it is impossible to estimate the gradient of reward functions without a complete generation process. Finally, we wish to highlight that this alignment process can run on only 1 consumer-level GPU, with memory usage less than 15GB. In contrast, current tuning-based methods require much more computing resources, typically 8 advanced GPUs.
@@
Table 1: Performance comparison. For SD v1.5 and PRNO, we annotate the generation time below the name. For LGD, we annotate the number of samples used for Monte Carlo approximation. For SPIN, DDPO, and AlignProp, we annotate the estimated time for fine-tuning. All time are measured with respect to the GPU time on an A800.
@@ 5.3 Optimization for Non-Differentiable Reward Functions
Setting. In this section, we aim to explore the three extensions proposed in Section 4 for handling non-differentiable reward functions. For Method 1, we use ZO-SGD (Nesterov and Spokoiny, 2017), and compare it with our proposed Hybrid-1 and Hybrid-2. We consider using two reward functions: The Jpeg Compressibility score, which is the file size of the image after compressing it using the JPEG algorithm. This reward model was also used in (Black et al., 2023) and is intrinsically non-differentiable. The second reward function is the aesthetic score used in the previous section. We compare the three methods in terms of optimization steps and the final score of the reward function. For ease of demonstration, we do not add the probability regularization term to the optimization process. Moreover, it is important to note that in these three methods, the major time expenditure comes from estimating the gradient with finite samples. For a fair comparison among these methods, we set the number of samples for estimating the gradient separately so that the total time spent estimating the gradient is the same for all three methods.
@@
Results. The main results are visualized in Figure 8. Initially, we observe that the Hybrid-2 method is significantly faster than the other two methods, and the final score is also higher for both reward functions used in this experiment. Furthermore, in both scenarios, the ZO-SGD method exhibits the slowest optimization speed. Another interesting finding is the poor performance of Hybrid-1 in optimizing the Aesthetic Reward. This mainly occurs because the aesthetic reward model can only work with image inputs, which validates our initiative to propose the Hybrid-2 method in Section 4. Finally, when comparing the optimization of the aesthetic score in Figures 7 and 8, we can also note that using the true gradient results in substantially fewer steps than using the estimated gradient.

### v2

@@ 3 Out-Of-Distribution Reward-Hacking in Noise Optimization
It has been observed that when aligning generative models (e.g., including autoregressive language models or diffusion models) with reward functions, one can experience the so-called reward-hacking, i.e., the optimized samples yield high rewards but do not possess the desirable properties (Miao et al., 2024; Chen et al., 2024a). Generally, there are two different types of reward-hacking. In the first type, the reward model is undertrained, so the optimized samples score high rewards but remain barely distinguishable from the samples of the pretrained distribution. The second type is related to the generative model used – the optimized samples no longer fall within the support of the pretrained distribution after optimization. We denote this second type of reward-hacking as Out-Of-Distribution (OOD) Reward-Hacking. In this work, we will focus on this second type and reveal that OOD reward-hacking is a common issue in DNO.
We begin by presenting two examples using both a simple diffusion model from the Example 1 described above and the open-source image diffusion model SD v1.5 (Rombach et al., 2022). For the first example, we revisit the pretrained distribution displayed in Figure 4, modifying the reward function to $r(x)=-(x[1]-1.4)^{2}-(x[2]-1.4)^{2}$. Figure 6(a) exhibits the pretrained distribution of the diffusion models, where we note that every sample stays within the support. However, as illustrated in Figure 6(b), after 1000 gradient steps of optimization for the reward function, all the generated samples become out-of-distribution. In our second example, we examine optimization for the "whiteness" reward—specifically, the average value of all pixels in an image—using SD v1.5 as the image diffusion model. We start with the prompt "black duck", with the initial image depicted in Figure 6(c). After 50 gradient steps of optimization for the whiteness reward, it becomes apparent that the generated samples diverge from the original prompt "black duck" and transform into a white duck, as evidenced in Figure 6(d). Ideally, in the absence of reward-hacking, the generated samples should always adhere to the "black duck" prompt while incorporating the overall whiteness in the images.
@@
The Reasons Behind OOD Samples. In this work, one of our key contributions is to identify one critical cause of OOD reward-hacking in noise optimization. That is, the optimized noise vectors stray towards the low-probability regions of the high-dimensional standard Gaussian distribution; in other words, there is an extremely low chance of such noise vectors being sampled from the Gaussian distribution. As diffusion models are originally trained with Gaussian noise, when the noise vectors originate from these low-probability areas—such as vectors comprised entirely of zeros—the neural network within the diffusion models may incur significant approximation errors for these particular inputs. This error, in turn, prompts the generation of out-of-distribution samples. In the subsequent section, we introduce a novel method to measure the extent to which noise is part of the low-probability region by leveraging the classical concentration inequalities for high-dimensional Gaussian distributions.
@@ 3.1 Quantifying Low-Probability Region via Concentration Inequalities
An important point to note is that the standard Gaussian distribution is invariant to permutation, i.e., for any permutation matrix $\Pi$, if $z$ follows a standard Gaussian distribution, the permuted vector $\Pi z$ will have the same probability behavior. With this insight, to increase the robustness of the probability measure $p_{1}$ and $p_{2}$, a natural idea is to examine the probability of many permuted vectors. Specifically, given $q$ permutation matrices $\Pi_{1},...,\Pi_{q}$, we define the following indicator metric,
@@
Interpretation of $P(z)$. If the probability $P(z)$ is low, it implies that there exists a permutation matrix $\Pi_{i}$ such that the noise vector $\Pi_{i}z$ is in the low-probability region of the standard Gaussian distribution. Therefore, due to the permutation-invariant property, the noise vector $z$ is also less likely to be sampled from the standard Gaussian distribution. In practice, we utilize randomly generated permutation matrices and have found that setting $q=100$ results in empirically good performance. In Appendix C, we provide some visualized empirical evidence to show that $P(z)$ serves as a good indicator for determining if the generated samples are OOD.
@@ 4 Tackling Non-Differentiable Reward Functions
In the previous section, DNO method has been applied to optimize differentiable reward functions. However, in many applications the ground-truth gradient of the reward function is unavailable. Such non-differentiable properties can arise from various scenarios; here we present two representative cases. Firstly, the reward may be computed through simulation-based procedures, such as the JPEG-compressibility employed in DDPO (Black et al., 2023), which calculates the size of an image in bits after running a JPEG compression algorithm. Additionally, the reward function itself may be a black box provided through online API providers, as in the setting considered in (Sun et al., 2022). This scenario is common when the reward function is a large neural network model, like those in (Wang et al., 2024a; Lin et al., 2024), making it impossible to directly obtain the gradient for optimization.
To address the scenarios mentioned above, in this section, we explore adapting the noise optimization approach to handle the optimization of non-differentiable reward functions by estimating the gradient with function values. Specifically, we explore three methods under this setting.
@@
Method 2. It is worth noting that for the mapping $r\circ M_{\theta}(\cdot)$, only the gradient of the reward function $r(x)$ is not available, and we are still able to compute the gradient of $M_{\theta}(z)$, since the noise-to-sample mapping itself is intrinsically differentiable. Therefore, a straightforward idea is to adopt a hybrid gradient approach—only to estimate the gradient of $r$, while using the ground truth gradient for $M_{\theta}(z)$. Specifically, we denote that the initial noise is $z$, and the generated sample is $x=M_{\theta}(z)$. Firstly, we can estimate the gradient of $\nabla r(x)$ in a similar fashion with the ZO-SGD (Nesterov and Spokoiny, 2017):
@@
where $\mu$ is the coefficient for perturbation, and $C_{1}$ is some constant. With the estimated gradient $H_{1}(x)$ for the reward function $r(x)$, we can use the following estimated gradient $G_{1}(z)$ for optimization:
$G_{1}(z)\stackrel{{\scriptstyle\text{def.}}}{{=}}H_{1}(x)\cdot\nabla_{z}M_{\theta}(z)\approx C_{1}\nabla_{z}r\circ M_{\theta}(z),$ (19)
where the main idea is to replace the ground truth $\nabla r(x)$ in the chain-rule of differentiating $r(M_{\theta}(z))$. We refer this method as Hybrid-1 in the following sections.
Remark 3.
As one can observe, the computation of (19) involves the Jacobian $\nabla_{z}M_{\theta}(z)$. However, it is important to note that when we only require the vector-Jacobian product $H_{1}(x)\cdot\nabla_{z}M_{\theta}(z)$, it is unnecessary to compute the full Jacobian $\nabla_{z}M_{\theta}(z)$. In Appendix D, we describe an elegant and efficient way to implement (19) using an auto-differentiation technique.
@@
Following a similar proof in (Nesterov and Spokoiny, 2017), we can also show that $H_{2}(x)\approx C_{2}\nabla r(x)$ for some constants $C_{2}$. As we can see, when computing the gradient (20), we ensure that we query the reward model $r(\cdot)$ with only samples that are within the manifold of the pretrained distribution. Similar to Hybrid-1, we can adopt a gradient estimator $G_{2}(z)$ with $H_{2}(x)$:
$G_{2}(z)\stackrel{{\scriptstyle\text{def.}}}{{=}}H_{2}(x)\cdot\nabla_{z}M_{\theta}(z).$ (21)
We refer this method (21) as Hybrid-2. As we will see in Section 5.3, this Hybrid-2 method is significantly faster than the other two in terms of optimization speed.
@@ 5 Experiments
In this section, we aim to demonstrate the effectiveness of the method proposed above. For all subsequent experiments, we utilize Stable Diffusion v1.5 (Rombach et al., 2022) as the base model for noise optimization. For each figure, we perform the optimization using 1,000 different random seeds and report the average value along with the standard deviation (std) of the results. For comprehensive details regarding the implementation of our proposed methods, as well as information on hyperparameters, we refer readers to Appendices D and E. Additionally, we provide examples to visualize the optimization process in Appendix A. For all the following experiments, unless explicitly stated otherwise, a single run of DNO is performed on a single A800 GPU.
@@ 5.1 Examining the Effectiveness of Probability Regularization
Setting. In this section, we design an experiment to demonstrate the effectiveness of probability regularization proposed in Section 3.2. We consider two settings: The first involves optimizing the whiteness reward, i.e., the average value of all pixels—the larger this value is, the whiter the image becomes—with the prompt "black <animal>", where the token <animal> is randomly selected from a list of animals; The second setting involves optimizing the blackness reward, i.e., the negative of the whiteness reward, with the prompt "white <animal>".
The primary insight behind designing such experiments is the clear contradiction between the prompt and the reward, making it easier to trigger the OOD reward-hacking phenomenon. Moreover, it is straightforward to verify whether the generated samples is out-of-distribution by simply examining the color of the generated animals. In these experiments, we will compare the noise optimization process with and without probability regularization to understand the capability of probability regularization in preventing the OOD reward-hacking phenomenon.
Figure 7: Comparison of DNO with and without probability regularization. Upper row: Optimizing for the whiteness reward, i.e., the average value of all pixels in the images. Lower row: Optimizing for the blackness reward, i.e., the negative of the whiteness reward. The x-axis refers to the number of gradient ascent steps during optimization.
Measuring the Degree of OOD. As observed in this experiment, the reward function contradicts the input prompt, leading to inconsistency between the generated samples and the prompt. Therefore, we utilize the CLIP Score (CS) (Radford et al., 2021), a commonly used metric for measuring the semantic similarity between images and text descriptions, to gauge the degree of OOD for the generated samples. The higher the CS, the less likely the sample is to be out-of-distribution (OOD).
@@
Figure 8: Examples of optimized images with and without regularization. (a)(b): Prompt is "black rabbit"; (c)(d): Prompt is "white chicken".
Results. In Figure 7, we first observe that adding the regularization term leads to a mildly slower optimization process. However, the generated samples are much more consistent with the prompt throughout the entire optimization process, as reflected by the CS curve. The trajectory of $P(z)$ further corroborates that it is a good indicator of the OOD phenomenon, as it is positively associated with the CS. It is worth noting that the regularization effect provided by probability cannot be fully replicated by simply employing early-stopping in non-regularized optimization, as illustrated in Figure 7. We offer qualitative examples in Figure 8 to highlight the differences between optimized samples with and without regularization. Figure 8(a) and 8(b) are optimized images with whiteness reward and prompt "black rabbit". Figure 8(c) and 8(d) are optimized images with blackness reward and prompt "white chicken". In the two pairs of images, we select the optimized images that achieved the almost the same reward value from the optimizations with and without regularization, respectively. It is evident that at an equivalent level of reward, the non-regularized samples are more prone to be out-of-distribution, as the color of the generated animals is less likely to match the prompt.
@@ 5.2 Benchmarking Three Reward Models Trained on Human Feedback Data
Setting. In this section, we investigate the performance of the proposed method using three common reward models trained from human feedback data, specifically Aesthetic Score (Schuhmann et al., 2022b), HPS-v2 score (Wu et al., 2023), and PickScore (Kirstain et al., 2023), respectively. In this experiment, we also compare noise optimization with and without probability regularization. However, compared to the reward function used in the previous section, using these reward models presents a lesser chance for the optimized image to be OOD. In this case, to measure the benefit of probability regularization in maintaining the quality of the generated sample, we consider using the other two reward models as test metrics when optimizing one of them. For example, when optimizing the Aesthetic Score, we use the HPS Score and Pick Score as test metrics.
@@
Figure 9: Comparison of running DNO with three human-like reward models, with and without regularization. When optimizing one reward model, the other two are used as test metrics. A, H, P are short for Aesthetic Score, HPS Score, and Pick Score, respectively. The name for each line comprises the used reward model and whether the regularization is used. For example, A + w/ reg means optimizing aesthetic score with regularization.
Comparing to Existing Alignment Methods. We also summarize the performance of our proposed method in Table 1 and compare it to the major existing alignment methods discussed in the introduction. As we can see, the proposed method can easily outperform state-of-the-art tuning-based alignment methods without any fine-tuning on the network models, within a reasonable time budget for generation. We found that another tuning-free method, LGD (Song et al., 2023a), performs badly with these complex reward functions, since it is impossible to estimate the gradient of reward functions without a complete generation process. Finally, we would like to highlight that this alignment process can run with memory usage of less than 15GB, and thus can easily fit into just one consumer-level GPU. In contrast, current tuning-based methods require significantly more computing resources, typically 4-8 advanced GPUs like A100 (Black et al., 2023; Prabhudesai et al., 2023).
@@
Table 1: Performance comparison. For SD v1.5 and PRNO, we annotate the generation time below the name. For LGD, we annotate the number of samples used for Monte Carlo approximation. For SPIN, DDPO, and AlignProp, we annotate the estimated time for fine-tuning. All time costs in the table are measured with respect to the GPU time on a single A800 GPU. Baselines: LGD (Song et al., 2023a), SPIN (Yuan et al., 2024), DDPO (Black et al., 2023), AlignProp (Prabhudesai et al., 2023).
@@ 5.3 Optimization for Non-Differentiable Reward Functions
Setting. In this section, we aim to explore the three extensions proposed in Section 4 for handling non-differentiable reward functions. For Method 1, we use ZO-SGD (Nesterov and Spokoiny, 2017), and compare it with our proposed Hybrid-1 and Hybrid-2. We consider using two reward functions: The Jpeg Compressibility score, which is the file size of the image after compressing it using the JPEG algorithm. This reward model was also used in (Black et al., 2023) and is intrinsically non-differentiable. The second reward function is the Aesthetic Score used in the previous section. Unlike previous experiments, in this section we treat the Aesthetic Score as a non-differentiable reward function for optimization. The goal is to simulate a scenario where the neural network model of the Aesthetic Score can only be queried via an API provider that returns the score rather than its gradient. We compare the three methods in terms of optimization steps and the final score of the reward function. For ease of demonstration, we do not add the probability regularization term to the optimization process. Moreover, it is important to note that in these three methods, the major time expenditure comes from estimating the gradient with finite samples. For a fair comparison among these methods, we set the number of samples for estimating the gradient separately so that the total time spent estimating the gradient is the same for all three methods.
@@
Results. The main results are visualized in Figure 10. Initially, we observe that the Hybrid-2 method is significantly faster than the other two methods, and the final score is also higher for both reward functions used in this experiment. Furthermore, in both scenarios, the ZO-SGD method exhibits the slowest optimization speed. Another interesting finding is the poor performance of Hybrid-1 in optimizing the Aesthetic Reward. This mainly occurs because the aesthetic reward model can only work with image inputs, which validates our initiative to propose the Hybrid-2 method in Section 4. Finally, when comparing the optimization of the aesthetic score in Figures 9 and 10, we can also note that using the true gradient results in substantially fewer steps than using the estimated gradient.